# The low dimensional homology groups of the elementary group of degree two

### v2

@@ 4. The first homology of $\mathit{{\rm E}}_{2}$
The second page of the spectral sequence $E_{\bullet,\bullet}$ is as follows:
[…]
Now by an easy analysis of the above spectral sequence, we obtain the exact sequence
@@
Let study the composite map $\bar{I}(A)\rightarrow I(A)\overset{d_{2,0}^{2}}{\longrightarrow}\mathcal{G}_{A}\oplus A_{A^{\times}}$. Take the element $\langle\!\langle a\rangle\!\rangle\in\bar{I}(A)$, $a\in{A^{\times}}$. This is represented by
@@
$[\ ]\otimes\partial_{2}((\boldsymbol{\infty},\boldsymbol{0},\boldsymbol{a})-(\boldsymbol{\infty},\boldsymbol{0},\boldsymbol{1}))\in B_{0}(\mathit{{\rm E}}_{2}(A))\otimes_{\mathit{{\rm E}}_{2}(A)}Z_{1}(A^{2}),$
which we denote it again by $X_{a}$. Thus $d^{2}_{2,0}(\langle\!\langle a\rangle\!\rangle)=d^{2}_{2,0}(\overline{X_{a}})$. To calculate the above composition we look at the diagram
@@
$[gh]\otimes(\boldsymbol{\infty})=([g]+[h])\otimes(\boldsymbol{\infty})\ \ \text{and}\ \ [h^{-1}]\otimes(\boldsymbol{\infty})=-[h]\otimes(\boldsymbol{\infty}).$
@@
$[D(a)^{-1}h^{-2}_{a^{-1}}h^{-1}_{a}h^{3}_{1}]\otimes(\boldsymbol{\infty})=\left[\left(\!\begin{array}[]{cc}\!\!a^{-1}&\!\!0\\ \!\!0&\!\!a\end{array}\!\!\!\right)\left(\!\begin{array}[]{cc}\!\!1&\!\!-2a-a^{-1}+3\\ \!\!0&\!\!1\end{array}\!\!\!\right)\right]\otimes(\boldsymbol{\infty}).$
@@
(i) Let $R$ be a local ring such that its residue field has at least four elements. Then there is $a\in R^{\times}$ such that $a^{2}-1\in R^{\times}$. Now for any $R$-algebra $A$ and for any $x\in A$, we have
@@
$H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\simeq\begin{cases}A/\mathfrak{m}_{A}^{2}&\text{if $|A/\mathfrak{m}_{A}|=2$}\\ A/\mathfrak{m}_{A}&\text{if $|A/\mathfrak{m}_{A}|=3.$}\\ \end{cases}$
(iii) A commutative semilocal ring $A$ is universal for ${\rm GE}_{2}$ if and only if none of the rings $\mathbb{Z}/2\times\mathbb{Z}/2$ and $\mathbb{Z}/2\times\mathbb{Z}/3$ is a direct factor of $A/J(A)$ [14, Theorem 2.14], where $J(A)$ is the Jacobson radical of $A$. Thus if $A$ is semilocal with the above condition, then by the above corollary
@@
(i) For a positive square free integer $d$, let $\mathcal{O}_{-d}$ be the ring of algebraic integers of $\mathbb{Q}(\sqrt{-d})$. Cohn proves that for $-d\neq-1,-2,-3,-7,-11$, $\mathcal{O}_{-d}$ is discretely normed [3, §6] and so is universal for ${\rm GE}_{2}$ [3, Theorem 5.2]. Since $\mathcal{O}_{-d}^{\times}=\{1,-1\}$, by the above corollary we have
@@
$H_{1}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-d}),\mathbb{Z})\simeq\begin{cases}\mathbb{Z}\oplus\mathbb{Z}\oplus\mathbb{Z}/6\oplus\mathbb{Z}/2&\text{if $d=5$}\\ \mathbb{Z}\oplus\mathbb{Z}\oplus\mathbb{Z}/6&\text{if $d=6$}\\ \mathbb{Z}\oplus\mathbb{Z}\oplus\mathbb{Z}/12&\text{if $d=15$}\\ \mathbb{Z}&\text{if $d=19$}\\ \end{cases}$
(see [22, Corollaries 11.2, 12.2, 15.2, 16.2]). Thus the natural map
@@
But it follows from [7, Theorem 1.4] that $H_{1}(\mathit{{\rm SL}}_{2}(k[X,Y]),\mathbb{Z})$ is very large.
@@ 5. The second homology of $\mathit{{\rm E}}_{2}$
We know that $\bar{I}(A)/\bar{I}^{2}(A)\simeq\mathcal{G}_{A}$ [23, Theorem 6.1.11]. Consider the commutative diagram with exact rows
00{0}${\bar{I}^{2}(A)}$${\bar{I}(A)}$${\mathcal{G}_{A}}$00{0}00{0}${I^{2}(A)}$${I(A)}$${\mathcal{G}_{A}\oplus A_{A^{\times}},}$
@@
[…]
@@
is the above mentioned map. For $z\in{A^{\times}}$, the element $\langle\!\langle z\rangle\!\rangle\in I(A)$ is represented by $[\ ]\otimes\partial_{2}(X_{z}-X_{1})$, where $X_{z}:=(\boldsymbol{\infty},\boldsymbol{0},\boldsymbol{z})$. Consider the diagram
@@
In [10, §7] Hutchinson has constructed a similar cycles as in above lemma over fields. Let $F$ be a field with at least four elements. Let $a,b\in F^{\times}$ and $\lambda\in F^{\times}$ such that $\lambda^{2}-1\in F^{\times}$. Then Hutchinson has constructed a cycle
@@
Let $A$ be a semilocal ring such that $2\in{A^{\times}}$ and all its residue fields have more that four elements. Then the group $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ is generated by the homology class of the cycles $([D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1})$, $([D(a)|D(b)]-[D(b)|D(a)])$, $a,b\in{A^{\times}}$, and $([E_{12}(x)|E_{12}(y)]-[E_{12}(y)|E_{12}(x)])$, $x,y\in A$:
@@
[…]
Let $X\in H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ and let $\theta(X)=\sum\langle\!\langle a_{i}\rangle\!\rangle\langle\!\langle b_{i}\rangle\!\rangle$. If $F(a,b):=([D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1})$, then $\theta(X-\sum\overline{F(a_{i},b_{i})})=0$. It follows from this that $X-\sum\overline{F(a_{i},b_{i})}$ sits in the image of $H_{2}(\mathrm{B}(A),\mathbb{Z})$. Consider the extension $1\rightarrow\mathrm{N}(A)\rightarrow\mathrm{B}(A)\rightarrow\mathrm{T}(A)\rightarrow 1$, where $\mathrm{N}(A):=\bigg{\{}\begin{pmatrix}1&b\\ 0&1\end{pmatrix}:a\in{A^{\times}},b\in A\bigg{\}}$. Since any residue field of $A$ has more than four elements, there is $\lambda\in{A^{\times}}$ such that $\lambda^{2}-1\in{A^{\times}}$. This implies that $A_{A^{\times}}=0$. It follows from this that for any $k\geq 0$, $H_{k}({A^{\times}},A)=0$ [16, Corollary 3.2]. By studying the Lyndon/Hochschild-Serre spectral sequence of this extension we obtain the exact sequence
@@ 6. The third homology of $\mathit{{\rm E}}_{2}$
$[a]-[b]+\langle a\rangle\bigg{[}\frac{b}{a}\bigg{]}-\langle a^{-1}-1\rangle\Bigg{[}\frac{1-a^{-1}}{1-b^{-1}}\Bigg{]}+\langle 1-a\rangle\Bigg{[}\frac{1-a}{1-b}\Bigg{]},$
@@
is a homomorphism of $\mathbb{Z}[\mathcal{G}_{A}]$-modules. Hutchinson defined the $\mathbb{Z}[\mathcal{G}_{A}]$-modules $\overline{\mathcal{RP}}_{1}(A)$ and ¯$\overline{\mathcal{RB}}(A)$ as follows
@@
¯$\overline{\mathcal{RB}}(A):=\ker(\bar{\lambda}_{2}|_{\overline{\mathcal{RP}}_{1}(A)}:\mathcal{RP}_{1}(A)\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}}))$
@@
If $Y_{\bullet}(A)\rightarrow\mathbb{Z}$ is exact in dimension $<4$, then the above map becomes an isomorphism.
@@
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathbb{Z}[\mathcal{G}_{A}]\overset{\bar{\partial}_{2}}{\rightarrow}E_{2,1}^{1}\rightarrow\mathcal{RP}_{1}(A)\rightarrow 0.$
@@
${\mu_{2}(A)\otimes_{\mathbb{Z}}\mathbb{Z}[\mathcal{G}_{A}]}$${E_{2,1}^{1}}$${\mathcal{RP}_{1}(A)\longrightarrow 0}$${\mu_{2}(A)}$${\mu_{2}(A).}$¯$\scriptstyle{\bar{\partial}_{2}}$$\scriptstyle{d_{2,1}^{1}}$
@@
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathcal{I}_{A}\rightarrow E_{2,1}^{2}\rightarrow\mathcal{RP}_{1}(A)\rightarrow 0.$
@@
Let $A$ be a semilocal ring. For any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$, let either $A/\mathfrak{m}$ be infinite or if $|A/\mathfrak{m}|=p^{d}$, then $(p-1)d>2(n+1)$ ($(p-1)d>2n$ when $A$ is a domain). Then $H_{k}({\rm T}(A),\mathbb{Z})\simeq H_{k}({\rm B}(A),\mathbb{Z})$ for all $k\leq n$.
@@
Now let $A$ be a general semilocal ring with the given condition. Then similar to the proof of [11, Proposition 3.19], we can show that $H_{n}(\mathrm{T}(A),k)\simeq H_{n}(\mathrm{B}(A),k)$, where $k$ is a prime field and $(p-1)d>2n$. Now the claim follows from [15, Lemma 2.3]. ∎
@@
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathcal{I}_{A}\rightarrow E_{2,1}^{2}\overset{d_{2,1}^{2}}{\longrightarrow}H_{2}(\mathrm{B}(A),\mathbb{Z})\simeq H_{2}(\mathrm{T}(A),\mathbb{Z})\simeq{A^{\times}}\wedge{A^{\times}}$
is given by $b\otimes\langle\!\langle a\rangle\!\rangle\mapsto b\wedge a$ (see [16, Lemma 4.1]). Thus we have the commutative diagram
@@
¯$\overline{\mathcal{RB}}(A)\rightarrow\mathcal{RB}(A).$
Now from the above commutative diagram we obtain a surjective map $E_{2,1}^{\infty}\twoheadrightarrow\mathcal{RB}(A)$. By an easy analysis of the main spectral sequence we obtain the surjective map
$H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow E_{2,1}^{\infty}.$
@@
Let $A$ be a semilocal ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. For any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$ let either $A/\mathfrak{m}$ be infinite or if $|A/\mathfrak{m}|=p^{d}$, then $(p-1)d>8$ ($(p-1)d>6$ when $A$ is a domain). If either $-1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
@@
See [18, Theorem 3.1, Theorem 3.2, Corollary 3.3] and use Proposition 6.1 and [18, Lemma 2.8]. ∎
@@
The connection between the third homology groups of $\mathit{{\rm E}}_{2}(A)$ and $\mathit{{\rm PE}}_{2}(A)$ is given in [19, Proposition 5.1].
@@
We may assume that ${\rm char}(F)\neq 2$. Since for any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$, $|A/\mathfrak{m}|\geq 4$, there is $a\in{A^{\times}}$ such that $a^{2}-1\in{A^{\times}}$. It follows from this that $M=A$. Thus by Theorem 4.1 we have $H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})=0$. Now by [19, Theorem 3.1] we obtain the desired exact sequence. ∎
@@
The main result of this section is as follows, which generalizes the main result of [17, Theorem 5.1, Corollary 5.2].
@@
Let $A$ be a semilocal ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. For any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$ let either $A/\mathfrak{m}$ be infinite or if $|A/\mathfrak{m}|=p^{d}$, then $(p-1)d>8$ ($(p-1)d>6$ when $A$ is a domain). If either $-1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
@@
where $(-1)\otimes\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)$ maps in $H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ to the class of the cycle which is obtained from the shuffle product of $[-1]\in B_{1}(\mu_{2}(A))_{\mu_{2}(A)}$ with the cycle $[D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1}\in B_{2}(\mathit{{\rm E}}_{2}(A))_{\mathit{{\rm E}}_{2}(A)}$.
@@
00{0}${\mathcal{K}}$${\displaystyle\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}}$${\mathcal{\mathcal{RB}}(A)}$00{0}00{0}${{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))}$${H_{3}(\mathit{{\rm PE}}_{2}(A),\mathbb{Z})}$${\mathcal{RB}(A)}$00{0}
@@
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\mathcal{K}\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow 0$
@@
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}\rightarrow\mathcal{RB}(A)\rightarrow 0$
@@
$D(a)\wedge D(b)$ maps to the homology class of the cycle $([D(a)|D(b)]-[D(b)|D(a)])$, $a,b\in{A^{\times}}$. Now from the commutative diagram

### v3

@@ 4. The first homology of $\mathit{{\rm E}}_{2}$
In this section, we show that the first homology group of $\mathit{{\rm E}}_{2}(A)$ fits into a natural exact sequence that almost completely describes its structure.
@@
By an easy analysis of the above second page of the spectral sequence $E_{\bullet,\bullet}$, we obtain the exact sequence
@@
Let’s study the composite map $\bar{I}(A)\rightarrow I(A)\overset{d_{2,0}^{2}}{\longrightarrow}\mathcal{G}_{A}\oplus A_{A^{\times}}$. Take the element $\langle\!\langle a\rangle\!\rangle\in\bar{I}(A)$, $a\in{A^{\times}}$. This is represented by
@@
[…]
which we denote again by $X_{a}$. Thus $d^{2}_{2,0}(\langle\!\langle a\rangle\!\rangle)=d^{2}_{2,0}(\overline{X_{a}})$. To calculate the above composition we look at the diagram
@@
$gh]\otimes(\boldsymbol{\infty})=([g]+[h])\otimes(\boldsymbol{\infty})\ \ \text{and}\ \ [h^{-1}]\otimes(\boldsymbol{\infty})=-[h]\otimes(\boldsymbol{\infty}).$
@@
$D(a)^{-1}h^{-2}_{a^{-1}}h^{-1}_{a}h^{3}_{1}]\otimes(\boldsymbol{\infty})=\left[\left(\!\begin{array}[]{cc}\!\!a^{-1}&\!\!0\\ \!\!0&\!\!a\end{array}\!\!\!\right)\left(\!\begin{array}[]{cc}\!\!1&\!\!-2a-a^{-1}+3\\ \!\!0&\!\!1\end{array}\!\!\!\right)\right]\otimes(\boldsymbol{\infty}).$
@@
(i) Let $R$ be a local ring such that its residue field has at least four elements. Then there is $a\in R^{\times}$ such that $a^{2}-1\in R^{\times}$. Now for any $R$-algebra AA and for any $x\in A$, we have
@@
[…]
(iii) A commutative semilocal ring $A$ is universal for ${\rm GE}_{2}$ if and only if none of the rings $\mathbb{Z}/2\times\mathbb{Z}/2$ and $\mathbb{Z}/2\times\mathbb{Z}/3$ is a direct factor of $A/J(A)$ [15, Theorem 2.14], where $J(A)$ is the Jacobson radical of $A$. Thus if $A$ is semilocal with the above condition, then by the above corollary
@@
(i) For a positive square free integer $d$, let $\mathcal{O}_{-d}$ be the ring of algebraic integers of $\mathbb{Q}(\sqrt{-d})$. Cohn proves that for $d\neq-1,-2,-3,-7,-11$, $\mathcal{O}_{-d}$ is discretely normed [3, §6] and so is universal for ${\rm GE}_{2}$ [3, Theorem 5.2]. Since $\mathcal{O}_{-d}^{\times}=\{1,-1\}$, by the above corollary we have
@@
[…]
(see [24, Corollaries 11.2, 12.2, 15.2, 16.2]). Thus the natural map
@@
But it follows from [7, Theorem 1.4] that $H_{1}(\mathit{{\rm SL}}_{2}(k[X,Y]),\mathbb{Z})$ is an infinite group.
@@ 5. The second homology of $\mathit{{\rm E}}_{2}$
In this section, we study the second homology group of $\mathit{{\rm E}}_{2}(A)$, where $A$ is a commutative ring such that $2\in{A^{\times}}$. The fundamental ideal $I(A)$ plays a central role in this context. More precisely, we investigate a surjective map from $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ onto $I^{2}(A)$.
@@
We know that $\bar{I}(A)/\bar{I}^{2}(A)\simeq\mathcal{G}_{A}$ [25, Theorem 6.1.11]. Consider the commutative diagram with exact rows
[…]
@@
$0\rightarrow Z_{1}^{Y}(A^{2})\rightarrow Y_{1}(A^{2})\rightarrow Z_{0}^{Y}(A^{2})\rightarrow 0,\ \ \ \ \ 0\rightarrow Z_{0}^{Y}(A^{2})\rightarrow Y_{0}(A^{2})\rightarrow\mathbb{Z}\rightarrow 0.$
@@
[…]
@@
In [10, §7] Hutchinson has constructed a similar cycles as in the above proposition over fields. Let $F$ be a field with at least four elements. Let $a,b\in F^{\times}$ and $\lambda\in F^{\times}$ such that $\lambda^{2}-1\in F^{\times}$. Then Hutchinson has constructed a cycle
@@
Let $A$ be a semilocal ring such that $2\in{A^{\times}}$ and all its residue fields have more that four elements. Then the group $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ is generated by the homology class of the cycles $([D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1})$, ([D(a)|D(b)][D(b)|D(a)])([D(a)|D(b)]-[D(b)|D(a)]), $a,b\in{A^{\times}}$, and $([E_{12}(x)|E_{12}(y)]-[E_{12}(y)|E_{12}(x)])$, $x,y\in A$:
@@
$H_{2}(\mathrm{B}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\overset{\theta}{\rightarrow}I^{2}(A)\rightarrow 0.$
Let $X\in H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ and let $\theta(X)=\sum\langle\!\langle a_{i}\rangle\!\rangle\langle\!\langle b_{i}\rangle\!\rangle$. If $F(a,b):=([D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1})$, then $\theta(X-\sum\overline{F(a_{i},b_{i})})=0$. It follows from this that $X-\sum\overline{F(a_{i},b_{i})}$ sits in the image of $H_{2}(\mathrm{B}(A),\mathbb{Z})$. Consider the extension $1\rightarrow\mathrm{N}(A)\rightarrow\mathrm{B}(A)\rightarrow\mathrm{T}(A)\rightarrow 1$, where $\mathrm{N}(A):=\bigg\{\begin{pmatrix}1&b\\ 0&1\end{pmatrix}:a\in{A^{\times}},b\in A\bigg\}$. Since any residue field of $A$ has more than four elements, there is $\lambda\in{A^{\times}}$ such that $\lambda^{2}-1\in{A^{\times}}$. This implies that $A_{A^{\times}}=0$. It follows from this that for any $k\geq 0$, $H_{k}({A^{\times}},A)=0$ [17, Corollary 3.2]. By studying the Lyndon/Hochschild-Serre spectral sequence of this extension we obtain the exact sequence
@@ 6. The third homology of $\mathit{{\rm E}}_{2}$
A potential refined Bloch–Wigner exact sequence is an exact sequence that describes the structure of the third homology group of $\mathit{{\rm E}}_{2}(A)$. Such an exact sequence has been established in a few cases [8], [11], [18]. We believe that in the present section we take a significant step toward a better understanding of such an exact sequence over semilocal rings.
@@
$a]-[b]+\langle a\rangle\bigg[\frac{b}{a}\bigg]-\langle a^{-1}-1\rangle\Bigg[\frac{1-a^{-1}}{1-b^{-1}}\Bigg]+\langle 1-a\rangle\Bigg[\frac{1-a}{1-b}\Bigg],$
@@
is a homomorphism of $\mathbb{Z}[\mathcal{G}_{A}]$-modules. Hutchinson defined the $\mathbb{Z}[\mathcal{G}_{A}]$-modules $\overline{\mathcal{RP}}_{1}(A)$ and $\overline{\mathcal{RB}}(A)$ as follows
@@
$\overline{\mathcal{RB}}(A):=\ker(\bar{\lambda}_{2}|_{\overline{\mathcal{RP}}_{1}(A)}:\mathcal{RP}_{1}(A)\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}}))$
@@
If $Y_{\bullet}(A)\rightarrow\mathbb{Z}$ is exact in dimension <4<4, then the above map becomes an isomorphism.
@@
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathbb{Z}[\mathcal{G}_{A}]\overset{\bar{\partial}_{2}}{\rightarrow}\mathrm{E}_{2,1}^{1}\rightarrow\mathcal{RP}_{1}(A)\rightarrow 0.$
@@
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathbb{Z}[\mathcal{G}_{A}]$$\mathrm{E}_{2,1}^{1}$$\mathcal{RP}_{1}(A)\longrightarrow 0$$\mu_{2}(A)$$\mu_{2}(A).$$\scriptstyle{\bar{\partial}_{2}}$$\scriptstyle{d_{2,1}^{1}}$
@@
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathcal{I}_{A}\rightarrow\mathrm{E}_{2,1}^{2}\rightarrow\mathcal{RP}_{1}(A)\rightarrow 0.$
@@
Let $A$ be a semilocal ring. For any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$, let either $A/\mathfrak{m}$ be infinite or if $|A/\mathfrak{m}|=p^{d}$, then (p1)d>2(n+1)(p-1)d>2(n+1) ((p1)d>2n(p-1)d>2n when $A$ is a domain). Then $H_{k}({\rm T}(A),\mathbb{Z})\simeq H_{k}({\rm B}(A),\mathbb{Z})$ for all $k\leq n$.
@@
Now let $A$ be a general semilocal ring with the given condition. Then similar to the proof of [11, Proposition 3.19], we can show that $H_{n}(\mathrm{T}(A),k)\simeq H_{n}(\mathrm{B}(A),k)$, where $k$ is a prime field and (p1)d>2n(p-1)d>2n. Now the claim follows from [16, Lemma 2.3]. ∎
@@
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathcal{I}_{A}\rightarrow\mathrm{E}_{2,1}^{2}\overset{d_{2,1}^{2}}{\longrightarrow}H_{2}(\mathrm{B}(A),\mathbb{Z})\simeq H_{2}(\mathrm{T}(A),\mathbb{Z})\simeq{A^{\times}}\wedge{A^{\times}}$
is given by $b\otimes\langle\!\langle a\rangle\!\rangle\mapsto b\wedge a$ (see [17, Lemma 4.1]). Thus we have the commutative diagram
@@
$\overline{\mathcal{RB}}(A)\rightarrow\mathcal{RB}(A).$
Now from the above commutative diagram we obtain a surjective map $\mathrm{E}_{2,1}^{\infty}\twoheadrightarrow\mathcal{RB}(A)$. By an easy analysis of the main spectral sequence we obtain the surjective map
$H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow\mathrm{E}_{2,1}^{\infty}.$
@@
Let $A$ be a semilocal ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. For any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$ let either $A/\mathfrak{m}$ be infinite or if $|A/\mathfrak{m}|=p^{d}$, then (p1)d>8(p-1)d>8 ((p1)d>6(p-1)d>6 when $A$ is a domain). If either $1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
@@
See [19, Theorem 3.1, Theorem 3.2, Corollary 3.3] and use Proposition 6.1 and [19, Lemma 2.8]. ∎
@@
The connection between the third homology groups of $\mathit{{\rm E}}_{2}(A)$ and $\mathit{{\rm PE}}_{2}(A)$ is given in [20, Proposition 5.1].
@@
We may assume that ${\rm char}(F)\neq 2$. Since for any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$, $|A/\mathfrak{m}|\geq 4$, there is $a\in{A^{\times}}$ such that $a^{2}-1\in{A^{\times}}$. It follows from this that $M=A$. Thus by Theorem 4.1 we have $H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})=0$. Now by [20, Theorem 3.1] we obtain the desired exact sequence. ∎
@@
The main result of this section is as follows, which generalizes the main result of [18, Theorem 5.1, Corollary 5.2].
@@
Let $A$ be a semilocal ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. For any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$ let either $A/\mathfrak{m}$ be infinite or if $|A/\mathfrak{m}|=p^{d}$, then (p1)d>8(p-1)d>8 ((p1)d>6(p-1)d>6 when $A$ is a domain). If either $1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
@@
where $(-1)\otimes\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)$ maps in $H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ to the class of the cycle which is obtained from the shuffle product of $1]\in B_{1}(\mu_{2}(A))_{\mu_{2}(A)}$ with the cycle $D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1}\in B_{2}(\mathit{{\rm E}}_{2}(A))_{\mathit{{\rm E}}_{2}(A)}$.
@@
[…]
@@
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\mathcal{K}\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow 0.$
@@
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}\rightarrow\mathcal{RB}(A)\rightarrow 0.$
@@
$D(a)\wedge D(b)$ maps to the homology class of the cycle ([D(a)|D(b)][D(b)|D(a)])([D(a)|D(b)]-[D(b)|D(a)]), $a,b\in{A^{\times}}$. Now from the commutative diagram